OpenStudy (anonymous):

What is the solution for x in the proportion 12 over 22 equals x over 99.? Describe your steps

OpenStudy (anonymous):

@Omar_Elboredy

OpenStudy (anonymous):

\[\frac{ 12 }{ 22 } = \frac{ x }{ 99 }\] will be equal to : \[\frac{ 12 }{ 22 } * 99 = 54\]